Fermi National Laboratory
 

Fermilab Shutdown Report for April 7, 2006 - April 14, 2006

 

LINAC

April 7 - 14

  • Running to NTF with the I- Ion Source
  • Beam Current 35 ma
  • Brought New Anode Supply On-line for the I- Haefely.
  • Low Energy RF PM is Complete.
  • High Energy RF PM Continued and is Complete.
  • Characterization of the Low Energy RF Systems Continues.
  • Troubleshooting LRF7 For the Reason It Tripped Substation 3.
  • LRF5 Modulator Problem Found and Repaired.
  • Found H- Haefely Problem and Repaired it.
  • Began Assembling Cables and Equipment to Begin Checking High Energy RF Phase and Amplitude Calibration

April 14 - 21

  • Continue Characterizing Low Energy RF Systems
  • Possibly Install Motor Drive for 10 MeV Collimator
  • Continue Test/Conditioning of Econco Switchtubes LRF7
  • Continue Checking HE RF Phase and Gradient Readback Calibration
  • Install New Digitizers for BPM Readbacks

Linac will begin turning on for 400 MeV beam after the power outages scheduled for April 22 - 24, 2006

Booster

  • MI-8 Extraction and Dump line under vacuum!
    • Next Week pull cables
  • ORBMP girder installed!
  • 400 MeV Line magnets placed/aligned
    • Connect up beam tube – start vacuum pump down
    • Move ORBMP PS to Booster gallery
  • Checked/Repaired water leaks
    • Test new CUB LCW pumps
  • RF repair/maintenance continues
  • Sump discharge re-route continues

Schedule Issues

  • MI-8 Dump on schedule – (tasks remaining)
    • Pull and terminate cables
    • Install kickers after HV testing
    • Septa connections and testing
  • ORBMP – connect and test PS
  • 400 MeV PS work on schedule
  • RF work – still lots to do
  • LCW and sump work – still some issues

Pbar

On Schedule

Main Injector

  • Continued installation of the new MI8 line multiwires.
  • Continued installation of MI8 collimator.
  • Completed Lambertson bakes at 608, 620.
  • Continued installation of a large aperture quad at 101.
  • Continued installation of large aperture quad at 402.
  • Began installation of large aperture quad at 521,522.
  • Installed the new loads on MI10 kickers.

Recycler

Damper Kickers

R32 line Vacuum Burst Investigation

Pelletron Work

  • New SF6 Skid
    • Shipped today
    • Arrive Tuesday/Wednesday of next week
  • Repair of the following Vacuum Components
    • IPA03 – power supply - Done
    • TPS S05
  • Replacement of shaft Bearings
  • Repair Collector Current read back- Done
  • Repair Pulser switch
  • Upgrade Fault protection system
  • Misc. Maintenance

Tevatron

A48 Collimator

B12 Quadrupole

TEL 1

Jobs Not Yet Started

  • Re align low beta quads at B0
  • As found A and E sectors with laser tracker
  • A 17 separator
  • TEL 1 installation

SwitchYard

They are ready to install the quadrupole at A0.

NuMI

  • Horn 2 Repair on Schedule for Next Tuesday and Wednesday
    • Added Shielding Blocks Placed by Work Cell per ALARA Plan
  • Target Chase Chiller Testing Started at Vendor
    • Expecting Return of Chiller by End of April
  • Target Hall Air Dehumidification – 5 x Desiccant Units
    • Requisition Cut with Prism for Ductwork
    • Electrical Work Continues to Progress
    • Exploring Additional Cable Paths thru Stripline Chase
  • Target Chase Fan Off by Noon Today. Tritium Testing Continues.
  • Target Profile Monitor Installation Complete in Pre-Target
    • Will Survey in on Monday. Still Problems with OTR.
  • Cleaning of Decay Pipe Passageway and Drains Complete
  • Level Runs in Target Hall Completing Today
  • Chase Condensate Tank Pump Replaced
  • Installation of New Drip Ceiling in Pre-Target Started

MiniBooNE

  • Horn
    • Initial testing April 14 (now).
    • Complete testing week of April 17.
  • Special LCW System
    • Week of April 24.
    • Require special tech coverage.
  • Alignment
    • Week of May 1
  • Conical Washers
  • Completed last week

Collider Detector at Fermilab

On schedule

D-Zero (D0)

  • Layer 0
    • Complete installation of junction card mounts
    • Push Run IIb beam pipe through Layer 0 support structure
    • Connect cooling for Layer 0
    • Begin installing junction cards/routing cables/testing connections
  • Level 1 Calorimeter Trigger
    • Continue connecting BLS cables to patch panels and test pulsing BLS cables
    • Continue firmware debugging/development/verification
  • Level 1 Central Track Trigger
    • Continue DFEA2 output verification

  • D-Zero attempts to minimize thermal cycles on the solenoid coils to limit risks to coil joints.
    • D sector controls apparently rebooted Friday April 7 without prior warning
    • Loss of discharge pressure causing Mycom compressors to unload
    • D-Zero is dependent upon Mycom compressors powered on Feeder 46b to recover helium, and benefits from advance notice regarding anticipated changes in status of power or controls

Mechanical

Work Continues

Cryogenics

Cryo Operations

  • Tevatron at LN2 temperature and LN2 usage is stable with vendor deliveries running only 2-3 truckloads per day due to the Tevatron being 1/4 at room temperature. Helium losses are steady at 15 kscf/day (1/2 of normal during TeV operations).
  • A1, A3, and D1 are at room temperature and isolated with LOTO.
  • B4, E2 and F4 are also at room temperature, have been reconnected by Cryo, and are currently under normal helium flow with anticipation of Division sign-off and cooldown to LN2 temperature soon. F4 house can be cooled down to LN2 temperature without the TEL-1 installed.
  • A3 has been completed by MSD, though spool piece conning towers may still be retrofitted (5 out of 8 leak) by Tech Division during this shutdown. Cryo has completed epoxy of helium leaks on Tevatron power leads, though they still need a snoop leak test.
  • Pbar D60 Debuncher vacuum work has been completed by Pbar Dept, thus Cryo personnel will restore helium/nitrogen connections and cool down tanks to LN2 temperature after Division sign-off.
  • ACNET application programs relating to FRIG operations continue to be tested on Linux consoles.
  • Nitrogen Reliquefier startup is delayed due to concern of “old” electrical cables being used for cross-connect between MSS and KRSS, especially due to high startup currents of 3000 hp TSI compressor. IR Booster compressor (400 hp) is running and purifying the plant.
  • MCR monitoring cryo operations.

CHL Maintenance:

  • General facility maintenance is in progress.
  • CHL Helium Compr “A” cooling tower installation has been completed with painting in progress. Some minor work is left, such as electrical/controls and fencing. (PICTURE below)
  • CHL Helium Compr “D” is being overhauled. (PICTURES below)
  • Helium Coldbox-2 controls upgrade is in progress.

FRIG Maintenance:

  • We have been finding some leaking quick-disconnect fittings on conning towers. We have none in stock and thus order is being placed to get spares.

Facility Engineering Services

  • There’s a Booster LCW shutdown on Tuesday for pump work.
  • Workers are cleaning the bus duct for the Master Substation.
  • The final portion of the zebra mussel cleaning is temporarily on hold.
  • There is ICW work on Casey’s Pond equipment.
  • There will be a CUB LCW pump test.

Instrumentation

Work continues.

RF

Work continues.

Controls

Work continues.

ES&H Upgrades

Safety System Upgrades

Keep discussing safety with your crews to keep our focus on getting our jobs done safely.

  • Data Acquisition System Complete
  • P-Bar Source
    • Verify collimator
  • MiniBooNE
    • Spare Horn is okay to test

EE Support

They will begin conditioning the A0 kickers next week.

Alignment

  • Tevatron
    • Completed Deep Rod installation.
    • Completed E0 Crystal Collimator
    • Completed IPM both vert and Horiz
    • Completed Flying wire cans at E-11.
    • Completed Flying wires cans at E17
    • Schedule Drops
    • Schedule As found two sectors.
  • Main Injector
    • Completed 402, and realignment of 620
    • Completed Deep Rod installation.
  • 8Gev
    • Completed New Texas multiwires will be installed in the MI8 line at 803, 826, 827, 829, and 836. These multiwires will need alignment.
    • Scheduled
  • P-Bar
    • Completed Set new AP2 beam stop on beam center once it is installed in the Transport enclosure.
    • Deep Rod installation.
  • 400Mev
    • Completed 400MeV Girder - Final survey
    • Scheduled 400 MeV - Final Survey
  • Booster
    • Completed AP2 beam stop
  • Recycler
    • No work Scheduled
  • MiniBooNE
    • Scheduled May 1
  • NuMI
    • Completed Level run between MI-60 to NuMI SR3 (Target Hall) approx. 1500 ft.
    • Completed NuMI Drops SR2, SR3 (Today)
    • Schedule
  • Add Ons
    • Tevatron
      • E17 resurvey (Check of as found by field crew)
    • Pbar
      • Completed Topo @ 8Gev area next week (Conduit installed)
      • Schedule Deep rod installation.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/14/2006    email Fermilab